OpenStudy (reneeshumpert):

Two number cubes are rolled. What is the probability that the sum of the numbers rolled is either a 3 or an 8? I hate these, need help . will fan and medal

OpenStudy (whpalmer4):

refer to the table I made you. how many ways to roll and 8? how many ways to roll a 3. add together, divide by 36.

OpenStudy (reneeshumpert):

my computer moving slow I was going to post the table but it keeps freezing

OpenStudy (anonymous):

possible events for a sum=12,21 No. of events=2 P(sum of 3)=2/36 possible events for a sum of 8=26,35,44,53,62 No. of events=5 P(sum of 8)=5/36 P(sum of 3 or 8) \[=\frac{ 2 }{ 36 }+\frac{ 5 }{ 36 }=?\]